OpenStudy (anonymous):

use the elimination method to solve the equations; 5u+3v=-15 2u+v=-7

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Regarding your problem, you'd first want to find some way to make a set of variables cancel when you add the equations together. In this case it would be easiest to multiply the entire second equation by -3 to allow the v to cancel with the 3v inthe other equation. Also it's important to remember to multiply the entire equation. anyway, you would end up with something like this: 5u+3v=-15 -6u-3v=21 you'd add the terms just like in a regular math addition problem, like so: -u=6 (remember that the v's cancel) From that, u would equal -6 and now you can put u back into any of the original equations to solve for b (I would suggest using the second equation for this since v is alone in that one)